Project

General

Profile

From 09-10-2014
Attending: Kyle, Fernanda, Jose, Gavin, Hayes, Michael, Mat

Hardware Watch:

- New limits are in place. (FD only noisy threshold moved higher)
- CSV Files are correct.
- Behind the scenes work on-going.

Far Detector:

*Cosmic trigger rate currently doubled and also on the 8th. 1 Hz incrementing by 2. (Contact Operations)
-PE spectrum still spiky (but not ADC) following calibrator change.

Jose
- DCM sync plots naming (correct now)
- Month plot are not being made
- Time stamp needs fixing
- Zero 3D track hits on the week sync plot

-FD DDT trigger are quite high

Near Detector:
PE dist looks odd?